### v1

# Deformable Radial Kernel Splatting

Abstract
Recently, Gaussian splatting has emerged as a robust technique for representing 3D scenes, enabling real-time rasterization and high-fidelity rendering. However, Gaussians’ inherent radial symmetry and smoothness constraints limit their ability to represent complex shapes, often requiring thousands of primitives to approximate detailed geometry. We introduce Deformable Radial Kernel (DRK), which extends Gaussian splatting into a more general and flexible framework. Through learnable radial bases with adjustable angles and scales, DRK efficiently models diverse shape primitives while enabling precise control over edge sharpness and boundary curvature. iven DRK’s planar nature, we further develop accurate ray-primitive intersection computation for depth sorting and introduce efficient kernel culling strategies for improved rasterization efficiency. Extensive experiments demonstrate that DRK outperforms existing methods in both representation efficiency and rendering quality, achieving state-of-the-art performance while dramatically reducing primitive count.
Figure 1: Gaussian Splatting vs. Our Deformable Radial Kernel (DRK) Splatting: Gaussian splatting requires thousands of Gaussians to approximate detailed textures and shapes. In contrast, our kernel efficiently fits the target pattern with just 30 primitives, achieving superior results.
[…]
Project page: https://yihua7.github.io/DRK-web/.
1 Introduction
3D Gaussian Splatting [24] (3D-GS) has emerged as a leading 3D representation method thanks to its rapid rasterization and superior rendering quality. Its key advantages– explicit point-based representation, flexibility for manipulation, and MLP-free rendering– have positioned it as a dominant approach in novel view synthesis. However, this success raises an intriguing question: Is the Gaussian kernel truly optimal for representing 3D scenes?
Natural scenes often consist of primitives with diverse shapes (see Fig. 2), such as rectangles, triangles, and ellipses, which cannot be fully captured by Gaussians alone. Moreover, the inherently smooth nature of Gaussian kernels makes them less effective at representing sharp transitions. As a result, densely packed or numerous Gaussians are frequently required to accurately model scene primitives with non-Gaussian geometries or sharp boundaries (see Fig. 1), leading to inefficiencies in both computation and memory.
Recent works have explored modifications to the Gaussian kernel to better handle discontinuities. For example, GES [18] adaptively adjusts the exponent values to control the sharpness of Gaussians, while still maintaining rotational symmetry, which limits its representational flexibility. DisC-GS [40] and 3D-HGS [28] introduce curve- and half-cutting techniques to the Gaussian kernel, respectively, improving its ability to capture discontinuous transitions. However, these cutting-based approaches remain fundamentally constrained by the inherent smoothness of the Gaussian kernel. Another line of research [17, 19, 14, 7] leverages 2D Gaussians to represent 3D scenes. While this approach offers better surface reconstruction compared to 3DGS, it is still constrained by the inherent limitations of the Gaussian framework, which restricts the flexibility of the representation.
In this work, our objective is to design new shape primitives that can adapt to complex 3D scenes. Motivated by the success of 2D Gaussian Splatting (2D-GS), we focus on developing adaptive 2D planar kernels. Specifically, we introduce the 2D Deformable Radial Kernel (DRK), a flexible planar primitive for representing 3D scenes (see Fig. 4). At its core, DRK extends traditional radial bases with learnable parameters that enable precise control over shape deformation through three key features: 1) multiple radial bases with learnable polar angles and scale parameters allow the kernel to adaptively stretch, rotate, and shrink, providing significantly more flexibility than fixed Gaussian functions; 2) a hybrid distance metric combining $L1$ and $L2$ norms with learnable weights enables DRK to naturally represent both curved surfaces and sharp geometric features; and 3) an adaptive piecewise linear remapping function provides fine-grained control over value distributions, particularly enhancing the representation of sharp boundaries. Despite its enhanced expressiveness, DRK maintains a simple parametric form that is easy to optimize and efficient to render. The DRK determines the location and orientation of the tangent plane through its center and rotation parameters, building upon these radial bases as its structural foundation to conform to target shapes, bypassing the geometric and smoothness constraints typically imposed by Gaussian functions. As shown in Fig. 2, a single DRK primitive can flexibly model a wide range of shape primitives, whereas many Gaussians are often required to represent similar shapes. The formulation is also backwards-compatible – DRK naturally reduces to a Gaussian kernel when using two orthogonal major bases with different decay rates, while achieving superior rendering quality with comparable computational efficiency.
We experimentally validate DRK on our proposed dataset (DiverseScene) covering various scenarios, including rich textures, fine geometry, specular effects, and large spatial scale. We also evaluate DRK on unbounded Mip-NeRF 360 [2] datasets. Experimental results demonstrate our method has advantages in preserving visual details more effectively and efficiently.
Overall, our work makes the following contributions:
  • We introduce Deformable Radial Kernel (DRK), a novel primitive that generalizes conventional Gaussian splatting. DRK enables flexible shape modeling through learnable radial bases, hybrid distance metrics, and adaptive sharpness control, significantly reducing the number of primitives needed for high-quality scene representation.
  • We develop a differentiable and efficient rendering pipeline for DRK, featuring polygon-based tile culling and cache-sorting strategies to enhance both rendering efficiency and view consistency.
  • We created DiverseScenes, a new dataset covering rich textures, fine geometry, specular effects, and large scenes for more effectively evaluating the ability to model complex scenarios of different algorithms.
2 Related Work
2.1 Novel View Synthesis
Novel view synthesis from multi-view images represents a fundamental challenge in computer graphics and computer vision. The field has evolved through increasingly sophisticated 3D scene representations. Traditional light field approaches [16, 8, 27] attempt to model the entire scene through a single 4D function of ray-slice intersections, but this implicit representation struggles with sparse inputs and complex geometry. Advancing beyond single-function modeling, Multi-Plane-Images (MPI) [48, 52, 55, 64, 35] introduced a layered representation using multiple depth planes, offering better 3D structure modeling. However, its discrete depth layering limits the handling of large viewpoint changes, particularly in rotational views. To achieve continuous depth modeling, subsequent works adopted explicit 3D proxies including meshes [53, 5, 9, 56], point clouds [38, 54, 34, 1], and voxels [45, 12, 30, 46]. These representations offer more flexible geometry modeling, with point-based methods further enhanced by neural networks [42, 43] for improved stability.
Most recently, continuous neural representations have emerged as a powerful alternative, replacing discrete geometric proxies with MLPs that directly map 3D coordinates to scene properties. NeRF [36] pioneered this approach with radiance fields, spawning numerous extensions for faster inference [13, 23, 37], large-scale scenes [61, 2, 50, 15], dynamic content [29, 39, 47], reflection modeling [3, 57, 26, 21], and stylization [62, 20, 11].
2.2 Gaussian Splatting
Point-based rendering approaches have made the rendering process of point clouds differentiable, enabling end-to-end training for novel view synthesis. 3D Gaussian Splatting (3D-GS) [24] introduced scene modeling using 3D Gaussian distributions, compositing the appearance of intersected Gaussians along viewing rays through $\alpha$-blending. This seminal work has inspired numerous extensions across various applications, including dynamic scene view synthesis [32, 60, 22], 3D content generation [65, 51], geometry reconstruction [33, 19], video representation [49] and neural network-enhanced high-fidelity view synthesis [59, 31]. The flexible representation, computational efficiency, and superior rendering quality of 3D-GS have established it as a prominent approach in the field.
The key innovation of 3D-GS lies in its use of Gaussian kernels for point effect decay, where the continuous nature of these functions facilitates the optimization of discrete representations. Subsequent research has focused on both constraining and extending these Gaussian kernels for specific applications. Notable developments include the constraint of 3D Gaussians to 2D space [17, 19, 7], yielding planar representations that better approximate scene surfaces for precise geometry reconstruction. Other works have extended the framework to 4D space [10, 58], enabling temporal dynamic modeling through 3D slicing of 4D Gaussians for view synthesis. To address limitations in representing sharp edges, researchers have proposed modifications to the Gaussian shape using curves [40] or hemispheric truncation [28].
3 Preliminaries
In this section, we first review 3D Gaussian Splatting (3D-GS) and 2D Gaussian Splatting (2D-GS), and their mathematical formulations (Sec. 3.1 and Sec. 3.2). We then introduce our general kernel splatting framework that extends beyond traditional Gaussian primitives (Sec. 3.3). Finally, we analyze the inherent limitations of Gaussian-based representations, motivating the need for our proposed deformable radial kernel formulation (Sec. 3.4).
3.1 3D Gaussian Splatting (3D-GS)
3D Gaussian Splatting represents scenes using colored 3D Gaussians [24]. Each Gaussian $G$ is parameterized by a 3D center $\mu$, covariance matrix $\Sigma=RSS^{T}R^{T}$ (where $R$ is a rotation matrix from quaternion $q\in\mathbf{SO}(3)$ and $S$ is a scaling matrix from vector $s$), opacity $o$, and spherical harmonic coefficients $sh$ for view-dependent appearance. A scene is represented as $\mathcal{G}=\{G_{j}:\mu_{j},q_{j},s_{j},o_{j},sh_{j}\}$.
Rendering involves projecting Gaussians onto the image plane with 2D covariance $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$ and center $\mu^{\prime}=JW\mu$. The pixel color $C(u)$ is computed via neural point-based $\alpha$-blending:
$\vspace{-2mm}\small C({u})=\sum_{i\in N}T_{i}\alpha_{i}\mathcal{SH}(sh_{i},v_{i}),\text{ where }T_{i}=\Pi_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $\mathcal{SH}$ evaluates spherical harmonics with view direction $v_{i}$, and $\alpha_{i}$ is:
$\small\alpha_{i}=o_{i}e^{-\frac{1}{2}({p}-\mu_{i}^{\prime})^{T}\Sigma_{i}^{\prime}({p}-\mu_{i}^{\prime})}.$ (2)
By optimizing these Gaussian parameters and adaptively adjusting their density, 3D-GS achieves high-quality scene representation with real-time rendering capabilities.
3.2 2D Gaussian Splatting (2D-GS)
Recognizing that real-world scenes primarily consist of surface structures where 3D Gaussians naturally compress into planar formations, recent works [19, 7] proposed 2D Gaussian Splatting (2D-GS). The opacity $\alpha$ is computed using local coordinates $(u,v)$ on the tangent plane:
$\alpha=o\cdot\exp\left(-\frac{1}{2}\left(\frac{u^{2}}{s_{u}^{2}}+\frac{v^{2}}{s_{v}^{2}}\right)\right),$ (3)
where $s_{u}^{2}$ and $s_{v}^{2}$ denote the variances along the U and V axes. This surface-centric formulation aligns with the success of polygon meshes in 3D modeling, offering a more compact scene representation.
Figure 2: Comparison of 3D-GS versus a single DRK: DRK achieves superior geometric fidelity with just one primitive compared to multiple Gaussians. We visualize the contours of 3D-GS and DRK to better illustrate primitive count, scale, and position.
3.3 General Kernel Splatting
We extend conventional 2D Gaussians by introducing a more general planar kernel splatting formulation (Fig. 3). Given a ray from position $r_{o}\in\mathbb{R}^{3}$ with direction $r_{d}\in\mathbb{R}^{3}$ intersecting a tangent plane centered at $\mu$ with rotation $R\in SO(3)$, the intersection point is:
$\small i=r_{o}+\frac{(\mu-r_{o})^{T}R_{z}}{r_{d}^{T}R_{z}}r_{d},$ (4)
where $R_{z}$ is the plane normal. The local UV coordinates at this intersection are:
$\small\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}R_{x}^{T}\\ R_{y}^{T}\end{pmatrix}(i-\mu),$ (5)
where $R_{x}^{T}$ and $R_{y}^{T}$ are the first and second rows of $R$. The kernel splatting function then maps these UV coordinates to a density value $\alpha$, determining the primitive’s shape.
Figure 3: Illustration of general planar kernel splatting: UV coordinates of ray-plane intersections are mapped to density values via a kernel function that determines the primitive’s shape.
3.4 Discussions
While 3D-GS demonstrates remarkable efficacy, our analysis reveals several inherent limitations. First, screen-space projection of Gaussians yields 2D elliptical distributions with rotational symmetry, constraining their ability to approximate diverse primitive shapes. Second, their conic curve boundaries (derived by the $L2$ norm) challenge the representation of linear edges and intermediate edge forms. Third, Gaussian distributions inherently couple decay rate with spatial extent through the covariance matrix—sharp features necessitate narrow distributions, making it challenging to simultaneously capture abrupt transitions and extended spatial regions.
These constraints necessitate millions of fine-grained Gaussians to approximate arbitrary shapes, leading to over-parameterization while still failing to achieve perfect fidelity under practical constraints. As shown in Fig. 2, even elementary shapes like triangles and rectangles—which can be described concisely—require numerous Gaussians for approximation, exhibiting artifacts like interior inconsistencies and external floating points (last column). Our deformable radial kernel (DRK) addresses these limitations through radial basis functions with varying lengths $s_{k}$ and polar angles $\theta_{k}$, controlled by parameters $\eta$ and $\tau$ for contour curvature and sharpness (Fig. 4).
4 Deformable Radial Kernel Splatting
Deformable Radial Kernel (DRK) is a novel primitive for 3D scene representation characterized by a set of parameters $\Theta=\{\mu,q,s_{k},\theta_{k},\eta,\tau,o,sh\}$, where:
  • $\mu\in\mathbb{R}^{3}$, $q\in SO(3)$, $o$, and $sh$ follow 3D-GS for position, orientation, opacity, and view-dependent appearance
  • $\{s_{k},\theta_{k}\}_{k=1}^{K}$ define radial basis lengths and angles that control the kernel shape
  • $\eta\in(0,1)$ and $\tau\in(-1,1)$ control boundary curvature and sharpness, respectively
In the following sections, we detail each component of DRK. First, we introduce the radial basis formulation (Sec. 4.1), showing that traditional 2D Gaussians are a special case within our framework. Then, we discuss $L1\&L2$ norm blending and edge sharpening in Sec. 4.2 and Sec. 4.3, which enable flexible control of boundary curvature and sharpness. Next, we explore efficient rasterization strategies in Sec. 4.4. Finally, we explain our optimization process in Sec. 4.5.
Figure 4: DRK defines a deformable shape using radial basis functions characterized by lengths $s_{k}$ and polar angles $\theta_{k}$, with parameters $\eta$ and $\tau$ governing the shape’s curvature and sharpness respectively.
4.1 Radial Basis
Our radial basis defines a kernel’s shape through $K$ control points in polar coordinates $\{s_{k},\theta_{k}\}$, where $s_{k}>0$ represents each radial length and angles are ordered as $0<\theta_{0}<...<\theta_{K}\leq 2\pi$. For any point with local tangent coordinates $(u,v)$, we compute its polar representation using $L2$ norm $r_{2}=\sqrt{u^{2}+v^{2}}$ and angle $\theta=\arccos(u/r_{2})$. Given a point lying between basis $k$ and $k+1$ (i.e., $\theta_{k}<\theta\leq\theta_{k+1}$), we define its relative angular distance as $\Delta\theta_{k}=\frac{(\theta-\theta_{k})\pi}{\theta_{k+1}-\theta_{k}}$. The kernel function is then formulated as:
$\small\alpha=o\cdot\exp\left(-\frac{r_{2}^{2}}{2}\left(\frac{1+\cos(\Delta\theta_{k})}{2s_{k}^{2}}+\frac{1-\cos(\Delta\theta_{k})}{2s_{k+1}^{2}}\right)\right).$ (6)
This formulation smoothly interpolates between adjacent radial bases using cosine weighting terms, ensuring continuous transitions in the kernel’s shape.
2D Gaussian as a Special Case.
Our formulation generalizes 2D Gaussian kernels. By setting $K=4$ with $\theta_{k}=\frac{k\pi}{2}$ and $s_{0}=s_{2}=s_{u}$, $s_{1}=s_{3}=s_{v}$, we can derive the 2D Gaussian form in Eq. (3). For points in the first UV quadrant where $\Delta\theta_{k}=2\theta$:
$\small\alpha$ $=o\cdot\exp\left(-\frac{r_{2}^{2}}{2}\left(\frac{\cos^{2}\theta}{s_{0}^{2}}+\frac{\cos^{2}\theta}{s_{1}^{2}}\right)\right)$
$=o\cdot\exp\left(-\frac{1}{2}\left(\frac{u^{2}}{s_{0}^{2}}+\frac{v^{2}}{s_{1}^{2}}\right)\right),$
where we use the identity $\cos(\Delta\theta_{k})=\cos(2\theta)=2\cos^{2}\theta-1$. The derivation extends to other quadrants by symmetry.
4.2 $L1\&L2$ Norm Blending
The scaling term $\frac{1}{\bar{s}^{2}}=\left(\frac{1+\cos(\Delta\theta_{k})}{2s_{k}^{2}}+\frac{1-\cos(\Delta\theta_{k})}{2s_{k+1}^{2}}\right)$ in Eq. (6) enforces smooth contours at radial endpoints, where $\frac{d\bar{s}}{d\Delta\theta_{k}}=0$ at $\Delta\theta_{k}=0$ or $\pi$. However, since radials converge at the kernel center and are not parallel, this $L2$ formulation inherently produces conic curves, limiting its ability to represent straight edges common in man-made environments.
To address this limitation, we incorporate $L1$ norm into our DRK formulation. For a point $(u,v)$ between radial endpoints $e_{i}=(s_{i}\cos(\theta_{i}),s_{i}\sin(\theta_{i}))$ and $e_{i+1}$, we compute its $L1$ norm as:
$\small r_{1}=\left\|\begin{pmatrix}e_{i}&e_{i+1}\end{pmatrix}^{-1}\begin{pmatrix}u\\ v\end{pmatrix}\right\|_{1}.$ (7)
This transformation maps the diamond-shaped $L1$ unit ball ($|x|+|y|=1,0\leq x,y\leq 1$) to the segment between adjacent endpoints, enabling straight-line boundaries.
We then introduce a blending weight $\eta\in(0,1)$ to smoothly interpolate between $L1$ and $L2$ norms, yielding our complete kernel function:
$\small\alpha=o\cdot\exp\left(-\frac{1}{2}\left(\eta r_{1}^{2}+(1-\eta)\frac{r_{2}^{2}}{\bar{s}^{2}}\right)\right).$ (8)
Note that $r_{1}^{2}$ requires no explicit scaling term as it is inherently scaled through the inverse transformation in Eq. (7). Please see Fig. 4 (right) for illustration.
4.3 Edge Sharpening
Figure 5: Sharpening function illustration.
As shown in Fig. 2, Gaussian functions inherently couple scale and edge sharpness through their variance parameter, making it challenging to achieve precise shape approximation even with thousands of Gaussians. To decouple these properties, we introduce a sharpening coefficient $\tau\in(-1,1)$ that adaptively adjusts edge transitions. Given the exponential term $g$ (i.e., the density value before opacity scaling) in Eq. (8), our sharpening function is:
$\small\Psi(g)=\left\{\begin{array}[]{llcl}\frac{1-\tau}{1+\tau}g&\text{if }&0&\leq g<\frac{1+\tau}{4},\\ \frac{1+\tau}{1-\tau}g-\frac{\tau}{1-\tau}&\text{if }&\frac{1+\tau}{4}&\leq g<\frac{3-\tau}{4},\\ \frac{1-\tau}{1+\tau}g+\frac{2\tau}{1+\tau}&\text{if }&\frac{3-\tau}{4}&\leq g\leq 1.\end{array}\right.$ (9)
As illustrated in Fig. 5, this piecewise function reshapes the alpha values toward 0 or 1, creating sharper edge transitions while maintaining the kernel’s spatial extent. The final kernel opacity is computed as $\alpha=o\cdot\Psi(g)$.
4.4 Rasterization
Low-pass Filtering.
Camera-captured images represent discretely sampled signals from scene rays, inherently limiting the maximum recoverable frequency of 3D content. Without proper frequency constraints, optimization tends to generate small, floating primitives that overfit to individual training views. These primitives become too fine-grained (smaller than ray coverage) and only contribute to specific rays, resulting in poor generalization.
While 3D-GS leverages EWA splatting for closed-form Gaussian low-pass filtering, DRK requires an alternative approach. Following Botsch et al. [4], we approximate filtering by taking the maximum between the kernel function and a view-dependent low-pass filter. Since DRK represents planar primitives, its frequency should be bounded on its tangent plane while potentially unbounded when viewed orthogonally, so we scale the filter size by the cosine of the view angle:
$\small\tilde{\alpha}=\max(\alpha,o\cdot\text{exp}(-\frac{1}{2|r_{d}\cdot R_{z}|^{2}}(\frac{\Delta p_{w}^{2}}{s_{l}^{2}}+\frac{\Delta p_{h}^{2}}{s_{l}^{2}}))),$ (10)
where $\Delta p_{h}$ and $\Delta p_{w}$ are image space distances between the pixel and the projected kernel center, $r_{d}\cdot R_{z}$ accounts for view-dependent scaling, and $s_{l}$ controls the low-pass filter radius.
Methods Simple Texture Geometry Specular Large Average
PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM
2D-GS 42.10 .0209 .9969 42.59 .0624 .9834 29.90 .0513 .9700 27.18 .0570 .9550 27.80 .2500 .8388 33.92 .0881 .9514
3D-GS 39.51 .0311 .9941 44.57 .0628 .9968 29.14 .0552 .9658 27.16 .0758 .9523 32.19 .2054 .9017 34.41 .0861 .9621
3D-HGS 41.41 .0199 .9931 46.19 .0397 .9912 30.42 .0406 .9598 27.27 .0512 .9447 33.17 .1672 .8721 35.68 .0637 .9521
GES 40.95 .0266 .9955 45.23 .0579 .9977 30.36 .0468 .9704 27.04 .0632 .9550 31.67 .2074 .8987 35.05 .0804 .9634
DRK (S2) 40.34 .0283 .9946 47.67 .0537 .9976 30.68 .0477 .9709 27.07 .0704 .9570 31.57 .2090 .8985 35.03 .0823 .9637
DRK (S1) 42.24 .0168 .9974 47.86 .0410 .9988 31.67 .0394 .9788 27.77 .0631 .9603 33.78 .1685 .9196 36.62 .0668 .9701
DRK 43.46 .0105 .9985 49.08 .0396 .9990 31.93 .0366 .9803 28.19 .0592 .9615 35.28 .1348 .9372 37.58 .0564 .9752
Table 1: Rendering quality evaluation on DiverseScene datasets across various categories
2DGS 3DGS 3DHGS GES DRK(S2) DRK(S1) DRK Num($\downarrow$) 359K 336K 373K 330K 42 K 109K 260K MB($\downarrow$) 83.6 79.7 89.6 78.1 12.3 32.1 76.6 FPS($\uparrow$) 251.3 247.1 154.5 227.4 234.9 119.2 77.5
Table 2: Evaluation results of average primitive number (Num), model size (MB), and rendering speed (FPS) on DiverseScene.
Kernel Culling.
Our rendering pipeline, based on 3D-GS, performs image rendering using 2D thread grids on CUDA. The process begins by dividing images into 16×16 tiles, with each tile corresponding to potentially intersecting primitives (kernels or 3D Gaussians). While 3D-GS projects the major axis of a 3D Gaussian onto the image plane to create a square axis-aligned bounding box (AABB) for tile coverage, this method can be computationally inefficient for primitives that project as highly anisotropic or elongated shapes, as shown in Fig. 6. This inefficiency becomes particularly evident in DRK due to its diverse shape representations. To address this, we introduce a more precise radius calculation and an improved tile culling strategy.
Figure 6: Polygon-based tile culling using radial basis endpoints enables efficient rasterization by eliminating untouched tiles.
The effective range of DRK is governed by its radial basis, whose endpoints define the kernel’s boundary. Following the 3-$\sigma$ principle, we define the boundary radius $s^{c}_{k}$, but extend it to account for opacity $o$ and sharpness $\tau$ parameters. The calibrated radial length $s_{k}^{c}$ is defined as:
$s^{c}_{k}=s_{k}\sqrt{-\log\left(\Psi^{-1}(\frac{e^{-3^{2}}}{o})\right)}.$ (11)
The derivation details are provided in the Supplementary Material. Using this calibrated length, we calculate each radial basis endpoint $v_{k}$:
$v_{k}=\mu+s_{k}^{c}(\cos\theta_{k}R_{x}+\sin\theta_{k}R_{y}).$ (12)
These endpoints form a polygon that accurately approximates the kernel’s boundary. By projecting this polygon onto the image plane, we can efficiently identify and cull tiles that don’t intersect with the kernel, making the rasterization more efficient.
Figure 7: Our cache-sorting uses $r_{t}$ as keys in cache-sorting to rectify the sorting order while preventing popping artifacts.
Cache Sorting.
3D-GS sorts the Gaussian using the depth of the center $\mu$, which can lead to inconsistencies across multiple views. While Radl et al. [41] address this by estimating the depth of the Gaussian at the point of maximum density and applying a hierarchical sorting solution to mitigate popping artifacts caused by inconsistent sorting, we propose a simpler yet effective solution. In our approach, the distance $r_{t}$ from the camera to the intersection between the ray and DRK is calculated as:
$r_{t}=\frac{(\mu-r_{o})^{T}R_{z}}{r_{d}^{T}R_{z}}.$ (13)
During rendering, each thread processes a pixel using an sorted array to cache the $r_{t}$ values and kernel indices, as illustrated in Fig. 7. The array is kept sorted by performing insert sorting for each new-coming kernel. If the array is full, the kernel with the smallest $r_{t}$ is popped out for processing. A dynamic cursor is used to record the scanning. We maintain an array length of 8, which is sufficient for approximating a more accurate sort. Fig. 7 demonstrates our method’s effectiveness in handling multiple overlapping kernels with identical center depths, comparing rasterization results with and without cache-sorting, preventing popping artifacts while maintaining rendering efficiency.
Figure 8: Qualitative comparisons of DRK with state-of-the-art methods across various scenarios show that DRK effectively captures sharp texture and geometry boundaries.
4.5 Optimization
Parametrization.
DRK attributes are modeled using unconstrained learnable parameters in (,)($\infty$,$\infty$)( - ∞ , ∞ ), with appropriate activation functions to ensure valid ranges. We apply a sigmoid function to constrain opacity $o$ to $(0,1)$, an exponential function for scale activation $s_{k}$, and a normalization function for rotation quaternion $q$. The sharpness $\tau$ is bounded within $(-0.1,0.99)$ through a composite function combining sigmoid and linear remapping, while the $L1\&L2$ blending weight $\eta$ is activated using a sigmoid function. For the basis angle $\theta_{k}$, we employ a three-step activation process: first applying a sigmoid function, then adding a residual term $\frac{1}{K-1}$ to maintain minimum angular separation, applying cumulative summation to enforce monotonic increase, and finally normalizing to make $s_{K}=2\pi$. The residual term prevents basis polar angles from exceeding $\frac{\pi}{2}$, thereby avoiding degradation in the representation. We set $K=8$ to balance flexibility and memory efficiency.
Model Training.
Following 3D-GS, we optimize model parameters and dynamically adjust kernel density through an adaptive training process. For DRK-specific parameters - sharpness $\tau$, blending weight $\eta$, basis angles $\theta_{k}$, and scales $s_{k}$ we set a uniform learning rate of $5e^{-3}$ and decay them gradually to the rate $1e^{-2}$$\times$ at the end of training (35K steps). We implement three density control configurations through the 2D screen gradient densification threshold and opacity pruning threshold pairs: ($5e^{-4}$, $5e^{-2}$) for density comparable to 3D-GS, ($1e^{-3}$, $5e^{-2}$) for Sparse Level 1 (S1), and ($2e^{-3}$, $1e^{-1}$) for Sparse Level 2 (S2).
5 Experiment
Figure 9: Comparisons on the MipNeRF 360 dataset show our method achieves clearer details and high-fidelity rendering.
5.1 Datasets and Evaluation Metrics
To evaluate our method’s performance across diverse scenarios, we focused on four key aspects: texture complexity, geometric detail, view-dependent effects, and scene scale. We collected 10 representative 3D scenes from Sketchfab [^1] https://sketchfab.com/, with two scenes per category: simple objects, richly textured surfaces, intricate geometric details, specular materials, and large-scale environments. The dataset, namely DiverseScenes, serves as a benchmark to assess our representation method’s capabilities. Detailed information about DiverseScenes can be found in the supplementary materials. For the evaluation on unbounded real-world scenes, we utilized the Mip-NeRF360 dataset [2], which contains multi-view images captured in open environments. Camera poses were estimated using COLMAP [44]. We assessed performance using three metrics: Peak Signal-to-Noise Ratio (PSNR) for overall reconstruction accuracy, Structural Similarity Index (SSIM) for perceptual quality and Learned Perceptual Image Patch Similarity (LPIPS) [63] for human-perceived visual fidelity.
5.2 Quantitative Comparisons
DiverseScenes. We compare our method, DRK, against state-of-the-art Gaussian representations: 3D-GS [24], 2D-GS [19], 3D-HGS [28], and GES [18], using their official implementations. To ensure a fair comparison focused on kernel representation, we adjust the gradient threshold for densification in these methods, aligning their primitive count with that of 3D-GS (300K+). The results for each dataset category are shown in Table 1, demonstrating that our method achieves the best average performance. This is followed by our sparser version, DRK (S1). DRK (S2) also performs comparably to other state-of-the-art methods, with significantly fewer primitives and lower space occupancy, while maintaining comparable FPS, as detailed in Table 2. Although our method requires more computation for the kernel function, the rendering framerates remains efficient (FPS $>$ 50).
Mip-NeRF360 Datasets. We also tested DRK on Mip-NeRF360 datasets, known for accurate camera poses and unbounded scenes. Results are shown in Tab. 3. Our method generates slightly more kernels on this dataset, so we also adjusted the 3D-GS densification threshold for comparison, labeled 3D-GS(L). On unbounded scenes, DRK still excels in perceptual quality (LPIPS, SSIM), it shows less advantage in PSNR, especially with sparser kernels. We hypothesize that DRK may tend to overfit distant regions with limited supervision. To address this, we used MiVOS[6] to mask image sequences, training on original images but evaluating only the foreground. The results (M-PSNR) indicate our method’s strength in well-supervised central regions.
Methods PSNR($\uparrow$) M-PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) Num($\downarrow$) Size($\downarrow$) 2D-GS 26.32 32.01 .2987 .7617 928 K 216.0 M 3D-GS 26.48 32.23 .3045 .7537 811 K 191.3 M 3D-GS (L) 26.94 32.75 .2688 .7799 1191 K 281.6 M 3D-HGS 26.74 32.51 .2999 .7561 842 K 202.4 M GES 26.62 32.11 .3047 .7542 824 K 195.0 M DRK (S2) 26.20 32.31 .2781 .7601 388 K 125.9 M DRK (S1) 26.40 32.56 .2601 .7722 551 K 161.9 M DRK 26.76 32.81 .2364 .7871 952 K 279.1 M
Table 3: Quantitative evaluation on Mip-NeRF360 scenes [2].
Methods $\tau=0$ $\eta=0$ $K=3$ $K=5$ $K=8$
DRK (S2) 34.97 34.82 33.93 34.47 35.03
DRK (S1) 36.22 36.04 35.50 35.88 36.62
DRK 37.46 37.27 36.60 36.85 37.58
Table 4: Ablation study on the impact of $\tau$, $\eta$, and $K$.
5.3 Qualitative Comparison
We conducted qualitative comparisons to highlight the advantages of DRK over other state-of-the-art kernel representations. Fig.8 shows results across different scene categories, where our method achieves better texture and geometry fitting with a comparable or smaller number of kernels. Fig.9 demonstrates that DRK effectively handles fine geometries, such as flowers and leaves, and detailed textures, like tags and bark, outperforming other representations that may have used more primitives (see Tab. 3).
5.4 Ablation Study
To gain a comprehensive understanding of DRK, we perform ablation studies on its attributes to assess their impact. We remove the effects of sharpening and curvature learning by setting $\tau=0$ and $\eta=0$, respectively. We also examine the impact of the hyper-parameter $K$. Ablation results on DiverseScenes are shown in Tab. 4. We found that $K$ significantly influences performance, with $K=3$ causing the greatest drop. Additionally, $\eta$ is more important than $\tau$, though both enhance representation capability.
6 Conclusion
We present DRK, a novel primitive representation that generalizes and enhances Gaussian splatting. By incorporating learnable radial bases, adaptive shape control, and efficient rendering strategies, DRK effectively addresses the fundamental limitations of Gaussian kernels in representing diverse geometric features. Our experiments on both synthetic and real-world scenes demonstrate that DRK achieves superior rendering quality with significantly fewer primitives, while maintaining computational efficiency through carefully designed rasterization optimizations.
Limitations
Despite its effectiveness, DRK has several limitations. First, the increased per-primitive computation affects rendering speed compared to basic Gaussian splatting, though this is largely offset by the significant reduction in primitive count. Additionally, DRK may exhibit reduced performance in regions with limited supervision, particularly in distant areas of unbounded scenes. Future work could explore more robust optimization strategies to address the challenge.
Supplementary Material
Outline
In this supplementary file, we provide additional applications and potential usages of DRK, an introduction to the DiverseScenes dataset, implementation details, and further results that could not be included in the main paper due to space constraints. The content is organized as follows:
  • Sec. S1: Seamless conversion of mesh models to DRK representations, bridging millions of 3D assets with high-fidelity reconstructed scenes.
  • Sec. S2: Introduction to the DiverseScenes dataset.
  • Sec. S3: Additional experimental results on public datasets and an analysis of limitations.
  • Sec. S4: Detailed implementation of the DRK framework.
S1 Converting Mesh to DRK
Existing ”Splatting” methods [24, 19, 18, 28] fail to seamlessly integrate with traditional assets. In contrast, our DRK framework provides a versatile representation that allows for effortless conversion of triangle and polygon faces into DRK. This is accomplished by placing the DRK endpoints at the polygon vertices and adjusting the parameters $\tau$ and $\eta$ to control sharpness, with both $L1$ and $L2$ curvature set to 1. This process eliminates the need for training data rendering or model optimization. Consequently, rich 3D assets can be seamlessly transferred into DRK, bridging the gap between millions of traditional 3D assets and our fast, high-quality reconstruction representation. Figure S1 illustrates the potential applications of Mesh2DRK.
Figure S1: The conversion from Mesh to DRK demonstrates the efficiency of using DRK to incorporate traditional 3D assets into any DRK scene at low cost.
We present examples of converting 3D mesh assets to DRK within seconds, without the need for training data preparation or optimization, in Fig. S2. Rendered depth and normal images are also provided. In these examples, the DRK is kernel-wise colored and shaded over the base color using the normal and predefined illumination. In the future, by assigning UV attributes to DRK and rendering UV maps, UV textures can be fetched and deferred rendering performed with the rendered normal. This capability allows DRK to handle traditional assets and compose scenes reconstructed from real-world multi-views and man-made artistic 3D assets.
Figure S2: Examples of converting various 3D assets. The normal of DRK can be used for shading under illumination. UV texture mapping is applicable for DRK in future implementations.
It is worth noting that cache-sorting has minimal impact on reconstruction quality but plays a crucial role in the conversion from mesh to DRK. This is because Mesh2DRK produces a compact geometric representation, where each DRK kernel represents a relatively larger unit of a mesh face compared to those learned from multi-view images, which use smaller units to capture high-frequency appearance details. Cache-sorting ensures that the sorting order is nearly correct, resulting in satisfactory conversion outcomes.
S2 DiverseScenes Dataset
We collected 10 scenes from Sketchfab [^2] https://sketchfab.com/, encompassing a variety of 3D scene types. As illustrated in Fig. S3, the dataset includes scenes with simple geometry and textures (e.g., McCree, House), detailed textures (e.g., Newspaper, Book), fine geometry (e.g., PalmTree, Dress), and large scales (e.g., Minecraft, Street). The training set consists of 200 views, and the test set includes 30 views, sampled from the unit sphere. For Minecraft and Street, there are 230 training views, with the train and test views simulating a walking camera along specified paths.
Figure S3: Overview of DiverseScenes. The dataset is composed of five categories: simple, specular, detailed geometry, fine texture, and large scale.
Scene Geometry Texture Scale Material Coarse Fine Coarse Fine Small Large Diffuse Specular McCree × × × × House × × × × Book × × × × Newspaper × × × × Dress × × × × PalmTree × × × × Armor × × × × DiscoBall × × × × Street × × × × MineCraft × × ×
Table S1: Summary of scene attribute annotations.
We manually annotated the attributes of each scene, as shown in Table S1. Some attributes overlap; for example, Armor and Street also feature fine textures. To clarify the performance of methods on DiverseScenes, we present the per-scene results in Table S2. These results are also summarized by scene categories in the main paper.
Methods McCree House PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) 2D-GS 40.95 .0067 .9985 43.25 .0351 .9953 3D-GS 39.15 .0115 .9975 39.87 .0507 .9907 3D-HGS 40.64 .0062 .9965 42.17 .0335 .9897 GES 40.48 .0086 .9982 41.42 .0446 .9927 Ours (S2) 39.76 .0104 .9977 40.92 .0462 .9914 Ours (S1) 40.90 .0067 .9985 43.57 .0269 .9963 Ours 41.70 .0047 .9989 45.22 .0163 .9980 Methods Book Newspaper PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) 2D-GS 38.20 .1199 .9673 46.97 .0048 .9995 3D-GS 42.25 .1211 .9944 46.89 .0045 .9991 3D-HGS 43.48 .0775 .9842 48.89 .0019 .9981 GES 43.49 .1122 .9960 46.97 .0036 .9993 Ours (S2) 43.28 .1042 .9961 47.75 .0032 .9991 Ours (S1) 45.57 .0904 .9976 49.76 .0023 .9992 Ours 47.28 .0801 .9983 50.88 .0019 .9993 Methods Dress PalmTree PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) 2D-GS 26.04 .0367 .9807 33.76 .0659 .9593 3D-GS 25.45 .0378 .9782 32.82 .0725 .9533 3D-HGS 27.30 .0217 .9773 33.54 .0595 .9423 GES 27.31 .0254 .9859 33.41 .0682 .9549 Ours (S2) 27.56 .0362 .9770 33.80 .0592 .9648 Ours (S1) 29.19 .0225 .9899 34.14 .0562 .9677 Ours 29.61 .0196 .9917 34.25 .0536 .9688 Methods Armor Discoball PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) 2D-GS 32.65 .0500 .9742 21.70 .0639 .9357 3D-GS 31.51 .0687 .9684 22.81 .0828 .9361 3D-HGS 32.26 .0518 .9512 22.28 .0505 .9381 GES 31.98 .0603 .9713 22.09 .0661 .9386 Ours (S2) 32.01 .0592 .9736 22.13 .0816 .9404 Ours (S1) 33.27 .0430 .9791 22.26 .0832 .9414 Ours 34.07 .0346 .9814 22.31 .0838 .9415 Methods Minecraft Street PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) 2D-GS 23.15 .4119 .7262 32.45 .0881 .9514 3D-GS 25.81 .3516 .8116 38.56 .0591 .9917 3D-HGS 26.10 .3070 .7568 40.23 .0274 .9873 GES 25.83 .3554 .8072 37.51 .0594 .9902 Ours (S2) 25.01 .3584 .8019 38.12 .0596 .9950 Ours (S1) 25.98 .3057 .8410 41.58 .0312 .9981 Ours 26.84 .2512 .8755 43.72 .0183 .9989 Methods Average PSNR($\uparrow$) LPIPS($\downarrow$) SSIM($\uparrow$) Num($\downarrow$) Size($\downarrow$) FPS($\uparrow$) 2D-GS 33.92 .0881 .9514 359K 83.6M 251.3 3D-GS 34.41 .0861 .9621 346K 82.0M 247.1 3D-HGS 35.68 .0637 .9521 373K 89.6M 154.5 GES 35.05 .0804 .9634 330K 78.1M 227.4 Ours (S2) 35.03 .0823 .9637 42K 12.3M 234.9 Ours (S1) 36.62 .0668 .9701 109K 32.1M 119.2 Ours 37.58 .0564 .9752 260K 76.6M 77.5
Table S2: We show the PSNR, LPIPS, and SSIM metrics for novel view synthesis on DiverseScenes.
S3 More Experiments & Limitation Analysis
Evaluation on NeRF-Synthetic Scenes.
We conducted a quantitative evaluation on NeRF-Synthetic [36] scenes, reporting PSNR scores, primitive numbers, and model sizes in Table S3. Unlike DiverseScenes, the training cameras for NeRF-Synthetic scenes (except Ficus) are sampled from the upper hemisphere, and some test views fall outside the range covered by the training views. This setup partially assesses performance on view extrapolation. Our DRK method demonstrates superior performance. While the kernel number for DRK is slightly larger than for 3D-GS [24], DRK (S1) maintains a compact number and outperforms other methods in PSNR scores. DRK (S3) has a very small kernel number and model size, with an average PSNR still comparable to others.
Scene 2D-GS 3D-GS 3D-HGS GES DRK (S2) DRK (S1) DRK Chair 34.88 35.83 34.29 34.05 34.38 35.28 35.61 Drums 25.67 26.15 26.29 26.05 25.90 26.12 26.13 Ficus 35.80 34.87 35.45 35.27 35.56 36.27 36.50 Hotdog 36.89 37.72 37.54 37.13 37.19 37.84 38.17 Lego 34.82 35.78 33.92 33.73 33.90 35.38 36.25 Materials 30.14 30.00 29.88 29.74 29.38 30.14 30.48 Mic 34.38 35.36 36.58 35.73 35.17 35.70 36.00 Ship 31.09 30.80 31.10 30.94 30.84 31.28 31.42 Avg PSNR 32.96 33.32 33.13 32.83 32.79 33.50 33.82 Num 107K 131K 83K 73K 32K 75K 158K Size 25.0M 31.1M 20.0M 17.4M 9.6M 22.0M 46.6M
Table S3: PSNR scores, primitive numbers, and model sizes on NeRF-Synthetic [36] scenes (transposed).
Evaluation on Tank&Temple Scenes.
To assess our method’s performance on more challenging scenes with imperfect camera conditions due to dynamic objects and changing exposures, we evaluated the Tank&Temple [25] datasets. We used 9 scenes in total, including 8 intermediate scenes and the Truck scene. We report the PSNR scores, foreground-only PSNR scores (M-PSNR), primitive numbers, and model sizes in Table S4. The results indicate that DRK faces significant challenges with this dataset, likely due to higher camera error estimated by COLMAP [44]. The Tank&Temple datasets are captured in dynamic environments with moving pedestrians and changing exposure, making camera estimation more difficult than in MipNeRF-360 [2], where objects are primarily diffuse and free from view-dependent effects, transients, or significant sunlight exposure changes. To further investigate the robustness of DRK, we conducted evaluations with noisy camera data.
Methods 2D-GS 3D-GS 3D-HGS GES Ours (S2) Ours (S1) Ours PSNR 20.65 21.09 21.59 20.58 20.20 20.31 20.41 M-PSNR 26.50 26.92 27.58 26.56 26.36 26.41 26.37 Num 1168K 275K 267K 259K 173K 212K 383K Size 271.9M 65.4M 64.2M 61.7M 50.9M 62.4M 112.6M
Table S4: Quantitative evaluation on Tank&Temple scenes.
Robustness against Camera Noise
To evaluate the performance of DRK under varying levels of camera noise, we simulated camera noise with different standard deviations (Std). The PSNR scores on DiverseScenes with noisy cameras are reported in Table S5. We observed that the PSNR scores of DRK drop significantly as the camera noise increases, whereas the performance of 3D-GS degrades more smoothly and slightly.
Noise Std 3D-GS DRK DRK (S1) DRK (S2) 00 34.41 37.58 36.62 35.03 $31e-$ 3 33.44 31.59 31.19 30.88 $32.5e-$ 3 31.32 29.27 28.89 28.60 $35e-$ 3 29.37 27.85 27.44 27.14
Table S5: Average PSNR scores on DiverseScenes of 3D-GS and DRK under different levels of camera noise.
Fig. S4 shows the rendering results of 3D-GS and DRK trained with both accurate and noisy cameras. When trained with accurate cameras, DRK achieves higher-quality rendering with sharper and clearer appearances. However, even with very small camera noise, the performance of DRK deteriorates significantly, producing blurrier and more chaotic results compared to 3D-GS. In contrast, 3D-GS maintains the ability to model the coarse appearance of the scene under noisy conditions. These results demonstrate that DRK is less robust to camera noise, which may explain its performance drop on the Tank&Temple dataset.
Figure S4: Rendering results of 3D-GS and DRK trained on cameras with and without noise.
Figure S5: Sorting accuracy comparisons: We found that pre-sorting tile-kernel pairs based on the nearest distance, combined with cache-sorting, achieves the highest accuracy. Cache-sorting is sufficiently effective in correcting most sorting disorders.
Figure S6: The ”Tensor Graph” of DRK, showing the dependence between parameters to optimize, intermediate variables, and the final output ($\alpha$).
S4 Method Details
Tensor Graph of DRK
To provide a detailed overview of DRK, we present its ”Tensor Graph,” which illustrates the flow from the learnable leaf parameters through the intermediate variables, ultimately leading to the outputs. The graph is shown in Fig. S6. Blue arrows represent data dependencies, along which gradients are back-propagated in reverse during optimization.
Cache-Sorting
To clarify the cache-sorting algorithm, we briefly summarize the process in Algorithm 1. As discussed in StopThePop [41], the backward processing must also be adjusted to proceed from front to back to maintain consistency with forward rendering.
Input: A cache chain with limited size, a new DRK (with an index and depth)
Output: The index of textitDRK or a status code
1
2if the cache is empty then
3       if the new DRK is invalid then
4             return a finish code;
5
6      Initialize the cache with the new DRK;
7       return success;
8
9
10if the new DRK is invalid then
11       Pop the DRK from the head;
12       return the index of the popped DRK;
13
14
15if the cache is full then
16       Mark the head DRK to be popped;
17
18
19Determine where to insert the new DRK by scanning the cache, guided by the DRK’s depth;
20 Adjust the pointers in the cache to insert the new DRK at the correct position;
21
22if the cache was full then
23       Pop the oldest DRK;
24
25
26return the index of the popped DRK (or success if none was popped);
27
Algorithm 1 Cache-Sorting
We evaluate the effectiveness of cache-sorting using a cache length of 8. DRK kernels are randomly sampled from the space, and the depths of DRK intersections processed in a front-to-back order are visualized in Fig. S5. Additionally, we assess the performance using metrics such as accuracy, Kendall’s Tau, and MAE. Our results show that in the pre-sorting stage (kernel-tile sorting), sorting based on the nearest distance between the DRK and the tile achieves the highest sorting accuracy. Sorting based on the most centric approach closely follows in performance. Both methods provide notable improvements compared to cache-sorting alone. Presorting with the nearest distance is also better than the vanilla presorting strategy on DRK. For further details and a more in-depth discussion, we refer readers to the StopThePop [41] paper, a pioneering work in this field.